OpenStudy (moongazer):

If a 75-kg runner can accelerate from standing start to a speed of 10m/s in 3 seconds, ho much power is generated? I'll just check my answer, please show your solutions. :)

OpenStudy (anonymous):

1250

OpenStudy (moongazer):

y 1250???

OpenStudy (anonymous):

yap 1250

OpenStudy (fretje):

We can calculate the kinetic energy of the runner after 3 seconds: Ekin = 0.5m.v² = 3750 Joule Power = dW/dt (power is work done per unit of time) The runners initial kinetic energy was zero , therefore is Ekin = W P = dW/dt = Ekin/t = 3750J/3s = 1250 Watt. (this is not instantaneous power but mean power <P>) <P>= W/dt (dt is specified time interval) It seems that runners can generate so much power but only for a short time. Once the velocity is build up, the maintenance power needed is much less. Unless the runner crashes and gives all the power back to Mother earth.